### v2

# VLM-Pruner: Buffering for Spatial Sparsity in an Efficient VLM Centrifugal Token Pruning Paradigm

Abstract
Vision–language models (VLMs) excel at image understanding tasks, but the large number of visual tokens imposes significant computational costs, hindering deployment on mobile devices. Many pruning methods rely solely on token importance and thus overlook inter-token redundancy, retaining numerous duplicated tokens and wasting capacity. Although some redundancy-aware approaches have been proposed, they often ignore the spatial relationships among visual tokens. This can lead to overly sparse selections of retained tokens that fail to adequately cover the regions of target objects. To address these limitations, we propose VLM-Pruner, a training-free token pruning algorithm that explicitly balances redundancy and spatial sparsity. We introduce a centrifugal token pruning paradigm that enables near-to-far selection while prioritizing the preservation of fine-grained object details. Moreover, we design a Buffering for Spatial Sparsity (BSS) criterion that defers the selection of spatially distant tokens. We further adopt a parallel greedy strategy to conduct token selection efficiently. To mitigate information loss from pruning, we selectively fuse salient information from the discarded tokens into the retained ones. Comprehensive comparisons demonstrate that VLM-Pruner consistently outperforms strong baselines across five VLMs with an 88.9% pruning rate, while delivering an end-to-end inference speedup. The code is available at https://github.com/Casey-bit/VLMPruner.
1 Introduction
Figure 1: Comparisons between baselines and VLM-Pruner. Compared with importance-driven FastV and redundancy-reduction DART and DivPrune at pruning rates of 66.7%, 77.8%, and 88.9%, VLM-Pruner consistently outperforms them across five VLMs.
Figure 2: Comparisons of the actual pruning effects between baselines and VLM-Pruner. Visual question answering cases with correct (green) and incorrect (red) responses; numbers (from 1 to 64) denote selection order.
Building on the powerful reasoning capabilities of large language models (LLMs) [gpt4, vicuna, llama], vision–language models (VLMs) [llava, llava15, llavanext, llavavideo, qwen2vl] integrate a visual encoder (e.g., ViT [vit]) to introduce visual modalities. This design enables VLMs to excel in tasks such as image captioning, visual question answering, video understanding, and multimodal reasoning. This integration significantly broadens their range of practical applications. However, the visual modality inherently carries high information density, particularly in high-resolution images [llavanext] and multi-frame videos [llavavideo], which results in a number of visual tokens that can exceed textual tokens by hundreds or even thousands of times. When these abundant visual tokens are merely projected and concatenated with textual tokens before being fed into the decoder layers of LLMs [llava], where attention computation exhibits quadratic complexity, the overall efficiency is substantially degraded.
As discussed above, the excessive number of visual tokens not only reduces computational efficiency but also introduces significant redundancy: 1) Cross-modal redundancy, where only visual regions relevant to the textual query are necessary; and 2) Intra-modal redundancy, where tokens representing similar regions of the same object carry overlapping information. Recent studies have addressed these issues through training-free token pruning methods that select a representative subset of visual tokens, typically focusing on two strategies: 1) importance-based pruning [fastv, hired, pyramiddrop, sparsevlm], which leverages intra- or cross-modal attention scores to retain salient tokens; and 2) redundancy-based pruning [dart, divprune], which employs greedy algorithms to iteratively choose tokens with low similarity to those selected.
However, the above strategies still perform suboptimally and exhibit notable drawbacks: 1) The importance-driven strategies focus on similar local regions often leads to redundancy. Methods based on visual importance like FastV [fastv] and on visual-language importance like SparseVLM [sparsevlm] often perform no better than random token pruning. In some cases they even perform worse [dart], because importance driven criteria tend to preserve multiple similar local regions around the same focused object, causing substantial redundancy. Although methods such as FiCoCo [ficoco] and MustDrop [mustdrop] apply localized penalties to prevent redundancy from overly clustered tokens, as shown in the person of Case 2 in Fig. 2, moderate clustering can, in fact, better preserve local details.; 2) The redundancy-reduction strategies leads to dispersed token selections. Methods targeting visual redundancy like DivPrune [divprune] and those addressing cross-modal redundancy like DART [dart] generate token selections that are dispersed yet incomplete, failing to capture the fine-grained details of objects. As shown in Fig. 2, edge background regions have low similarity to the main object, so DivPrune and DART are prone to selecting edge background tokens and jumping erratically between background and foreground, resulting in a scattered selection.
Accordingly, we propose a training-free centrifugal token pruning paradigm, termed VLM-Pruner, that effectively balances token redundancy and local-detail completeness. VLM-Pruner follows a near-to-far token selection process, starting with pivot tokens, then prioritizing spatially adjacent tokens, and finally recovering from discarded tokens. This ensures an orderly sequence and avoids the chaotic, incomplete coverage from arbitrary selection. For example, in the first case of Fig. 2, four tokens closely aligned in order on the fork at the left of the image are selected when compared to other baselines; in the third case, VLM-Pruner captures comprehensive truck details, including the body, tires, and windshield. Specifically, a minimal pivot token set with maximum minimum distance [maxmin] is firstly identified to coarsely represent distinct subjects. Secondly, the Buffering for Spatial Sparsity (BSS) criterion leverages the minimum spatial distance between candidate tokens and the selected set to preferentially select spatially adjacent tokens with low redundancy, thereby ensuring an orderly selection process. Tokens are then greedily chosen in ascending similarity order, with parallel processing for acceleration. Finally, the remaining tokens are matched to the selected set based on maximum similarity and merged through Similarity-Weighted Aggregation (SWA) to recover the outermost information. In summary, our major contributions are as follows:
  • We propose VLM-Pruner, a training free centrifugal token pruning paradigm for efficient VLMs that balances token redundancy and local-detail completeness.
  • We adopt the Buffering for Spatial Sparsity criterion to enforce relatively ordered token pruning and effectively alleviate scattered token distribution.
  • We conduct extensive experiments on 13 benchmarks across 5 VLMs. As shown in Fig. 1, VLM-Pruner delivers state-of-the-art results, with gains that grow as the pruning ratio increases.
Figure 3: Centrifugal token pruning paradigm of VLM-Pruner. (a) Pipeline: In the $i$-th decoder layer of the LLM, VLM-Pruner follows a near-to-far selection order, (b) starting with pivot tokens, (c) gradually expanding outward from neighborhoods, and (d) ultimately recovering the outermost information from the discarded tokens via SWA. The similarity computed under BSS criterion makes candidate tokens spatially closer to selected ones more likely to be chosen. Color transition from green to red indicates decreasing selection probability. $C$ and $S$ denote candidate and selected tokens, respectively. After applying BSS, the closer candidate $C_{2}$ is prioritized over $C_{1}$.
2 Related Work
2.1 Efficient Vision-Language Models
Vision–Language Models (VLMs) [vit] align visual encoders with large language models to support captioning, VQA, grounding, and reasoning, with image-based [llava, llava15, llavanext, qwen2vl] and video-based [llavavideo] variants. A common recipe couples LLMs, which perform strong at text processing, with high-capacity visual encoders such as ViT [vit], CLIP-ViT [clipvit], or SigLIP [siglip] via cross-modal alignment.
Efficiency efforts aim to preserve multimodal capability while cutting compute, memory, and latency. (i) Connector-tuning freezes the backbones and trains lightweight adapters, e.g., BLIP-2 [blip], avoiding full end-to-end retraining. (ii) Content-adaptive budgeting allocates resolution and token budgets based on input saliency, as in Qwen2-VL [qwen2vl]. (iii) Distill-and-prune compression transfers knowledge [kd] and removes modality-aware redundancy. EfficientVLM [efficientvlm] retains 44.3% parameters and yields the inference speedup with minimal quality loss. Expert Merging [zhang2025expert] integrates capabilities of multiple VLMs.
2.2 Token Pruning
Token pruning reduces the quadratic cost of attention by compressing tokens at their source. ToMe [tome] pioneered progressive token merging for ViTs [vit], inspiring VLM pruning methods that are either training-required or training-free. Among training-required approaches, ATP-LLaVA [atpllava] learns layer-wise, instance-specific pruning ratios via MLPs, but adds training overhead. We focus on training-free pruning, which falls into importance-based and redundancy-based families.
Importance-driven methods keep tokens with high (self- or cross-modal) attention: FastV [fastv] uses self-attention on visual tokens; SparseVLM [sparsevlm] jointly prunes visual/text tokens via cross-modal attention; HiRED [hired] and MustDrop [mustdrop] favor tokens strongly attending to [CLS]; PDrop [pyramiddrop] retains those attentive to the last-instruction token. Redundancy-reduction methods prefer diversity: DART [dart] incrementally selects tokens with low similarity to text–visual anchors, while DivPrune [divprune] greedily preserves tokens least similar to those already kept.
In practice, importance-driven pruning can overselect semantically overlapping tokens, whereas redundancy-reduction selection may be unstable when similarity gaps between background and salient regions are large, yielding dispersed coverage and missing fine-grained details.
3 Methodology
3.1 Overview
As shown in Fig. 3, we present a training-free centrifugal pipeline for retaining the most informative image tokens in VLMs, while ensuring comprehensive coverage of dense local object details. Specifically, VLM-Pruner follows a three-stage paradigm: (i) initialize a small set of diverse pivots in feature space via a max-min rule; (ii) expand the retained set with a parallel greedy procedure regularized by a BSS criterion that prioritizes densifying local neighborhoods before expanding outward; and (iii) re-inject the information of discarded tokens into their most similar retained pivots through a similarity-weighted aggregation.
3.2 Buffering for Spatial Sparsity (BSS) Criterion
We perform VLM-Pruner on the second layer of the LLM decoder to enforce an ordered near-to-far selection process, thereby preventing the token distribution from becoming too scattered. Given that the visual feature map has a spatial size of $H\times W$, it results in $N=H\times W$ tokens. We denote the token keyskeys as $\mathbf{K}\in\mathbb{R}^{N\times d_{k}}$ to define the identity of features, and the hidden_stateshidden\_states as $\mathbf{H}\in\mathbb{R}^{N\times d}$ to compute similarities, respectively. The spatial coordinate of token $i\!\in\!\{0,\dots,N-1\}$ is $\mathbf{p}_{i}=(x_{i},y_{i})$ with
$x_{i}\;=\;i\bmod W,\quad y_{i}\;=\;\bigl\lfloor i/W\bigr\rfloor.$ (1)
The grid distance and its maximum can be defined as
$D^{(\mathrm{sp})}_{ij}\;=\;\bigl\lVert\mathbf{p}_{i}-\mathbf{p}_{j}\bigr\rVert_{2},\quad D_{\max}\;=\;\sqrt{H^{2}+W^{2}}.$ (2)
To reduce the cost of a full $N\times N$ similarity, we first screen the channels based on variance. We compute the per-channel variance of $\mathbf{H}$ across tokens, then keep the top-$q$ channels (with $q=256$). Afterward, we form the reduced features as $\tilde{\mathbf{H}}\,=\,\mathbf{H}_{:\,\!,\mathcal{I}}\in\mathbb{R}^{N\times q}.$ We use cosine similarity
$M_{ij}\;=\;\frac{\tilde{\mathbf{H}}_{i}^{\top}\tilde{\mathbf{H}}_{j}}{\lVert\tilde{\mathbf{H}}_{i}\rVert_{2}\,\lVert\tilde{\mathbf{H}}_{j}\rVert_{2}},$ (3)
and precompute the matrix $M\in\mathbb{R}^{N\times N}$ once.
Our central claim is that prioritizing spatial proximity during token selection facilitates a more complete reconstruction of fine-grained object details. In other words, the selection process naturally follows a centrifugal paradigm, starting from locality and gradually expanding outward.
We denote $\mathcal{V}=\{0,\dots,N-1\}$ as the set of all tokens. We denote $\mathcal{S}\subseteq\mathcal{V}$ as the retained set, where $|\mathcal{S}|=\mathcal{R}$. We also denote $\mathcal{C}=\mathcal{V}\setminus\mathcal{S}$ as the candidate set, where $|\mathcal{C}|=N$-$\mathcal{R}$. We pose selection as a maximum diversity problem $\mathcal{F}$ with a constraint of spatial sparsity:
$\max_{S\subseteq V}\;\;\mathcal{F}(S)\;=\;\max\!\!\sum_{i\subseteq\mathcal{C},j\subseteq S}\!\min\widetilde{M}_{ij},$ (4)
where $\widetilde{M}_{ij}$ is a BSS-modulated similarity
$\delta_{i}(S)\;$ $=\;\min_{j\in S}D^{(\mathrm{sp})}_{ij}$,
$\bar{\delta}_{i}(S)\;$ $=\;\delta_{i}(S)/D_{\max},$ (5)
$\widetilde{M}_{ij}\;$ $=\;M_{ij}\,\bigl(1+\lambda\,\bar{\delta}_{i}(S)\bigr).$
When $\lambda=0$, the objective $\mathcal{F}$ reduces to a standard monotone submodular function [cost], for which the greedy algorithm achieves the classical approximation guarantee [greedy]. When $\lambda>0$, as the iterations proceed, the number of tokens in the selected set $\mathcal{S}$ gradually increases, causing the remaining candidate tokens to lie closer to the already selected ones in spatial space. Consequently, the normalized spatial distance term $\bar{\delta}_{i}(S)$ in Eq. (5) continuously decreases with each iteration. This monotonic decay preserves the overall submodular structure of $\mathcal{F}$ across iterations and therefore still permits an efficient greedy optimization scheme. Meanwhile, for $\lambda>0$, tokens that are spatially farther from the already selected set are assigned a larger scaling coefficient in their similarity computation, making them less likely to be selected in subsequent iterations. As shown in Fig. 3, this adaptive weighting mechanism effectively realizes the BSS effect, ensuring that nearby tokens are prioritized first, thereby promoting dense local coverage before the selection expands outward.
Algorithm 1 Centrifugal Token Pruning Paradigm with BSS
0:  $\mathbf{K}\!\in\!\mathbb{R}^{N\times d_{k}}$, $\mathbf{H}\!\in\!\mathbb{R}^{N\times d}$, coordinates $\{\mathbf{p}_{i}\}$; target token count $R$, pivot count $\kappa$, kept channels $q$, BSS strength $\lambda$, threshold schedule $\tau^{(0)}$, $\Delta\tau$, batch size $B$.
1:  Keep the top-$q$ channels with the highest variance to form $\tilde{\mathbf{H}}$; build the similarity matrix $M$ by Eq. (3).
2:  Max–min initialize $\kappa$ pivots to get $\mathcal{S}_{\kappa}$ via (6).
3:  Precompute $D^{(\mathrm{sp})}_{ij}$ and $D_{\max}$.
4:  $t\!\leftarrow\!0$, $\tau\!\leftarrow\!\tau^{(0)}$, $\mathcal{S}\!\leftarrow\!\mathcal{S}_{\kappa}$, $j\in\mathcal{S}$.
5:  while $|\mathcal{S}|<R$ do
6:   $\mathcal{C}\!\leftarrow\!\{1,\dots,N\}\setminus\mathcal{S}$.
7:   For each $i\!\in\!\mathcal{C}$, compute $\bar{\delta}_{i}$ and form $\widetilde{M}_{ij}$ via Eq. (5).
8:   Compute $r_{i}$ by Eq. (7); sort $i$ in descending $r_{i}$.
9:   for batches of $B$ candidates $i$ do
9:    Recompute $\bar{\delta}_{i}$ and $\widetilde{M}_{ij}$ via Eq. (5).
10:    if $\max_{j\in\mathcal{S}}\widetilde{M}_{ij}<\tau$ then
11:     $\mathcal{S}\leftarrow\mathcal{S}\cup\{i\}$; break if $|\mathcal{S}|=R$.
12:    end if
13:   end for
14:   break if no token added this loop.
15:   $t\!\leftarrow\!t+1$, $\tau\!\leftarrow\!\tau+\Delta\tau$.
16:  end while
17:  $\mathcal{D}\!\leftarrow\!\{1,\dots,N\}\setminus\mathcal{S}$; build clusters $\{\mathcal{D}_{j}\}$ via $j^{\star}(u)$.
18:  Compute $\mathbf{M}_{j}$ and update $\mathbf{H}_{j}$ by Eq. (10).
19:  return Selected indices $\mathcal{S}$ and updated $\mathbf{H}$.
3.3 Overall Token Pruning Pipeline
We briefly summarize our proposed centrifugal token pruning method, termed VLM-Pruner, in Algorithm 1.
3.3.1 Stage 1: Pivot Initialization
This initialization yields widely separated pivots that coarsely cover different semantic regions.
We denote $\kappa$ as the number of initial pivots. The token keyskeys inherently summarize the semantic characteristics of each token [attention] and contain less redundant information due to their lower dimensionality. Therefore, we select $\kappa$ diverse pivots $\mathcal{S}_{\kappa}$ in the token keyskeys space using a max-min strategy:
$j_{1}$ $=\arg\max_{j\in V}\lVert\mathbf{K}_{j}\rVert_{1},$
$j_{t}$ $=\arg\max_{j\in\mathcal{C}}\min_{j^{\prime}\in\mathcal{S}_{t-1}}\bigl\lVert\mathbf{K}_{j}-\mathbf{K}_{j^{\prime}}\bigr\rVert_{2},\quad t=2,\dots,\kappa,$ (6)
$\mathcal{S}_{t}$ $\leftarrow\mathcal{S}_{t-1}\cup\{j_{t}\}.$
Finally, we define the selected set as $\mathcal{S}\leftarrow\mathcal{S}_{\kappa}$.
3.3.2 Stage 2: Greedy Selection with BSS Criterion
A purely redundancy-reduction greedy selection may over-disperse tokens. To address this, we introduce the BSS criterion, which enforces an ordered near-to-far selection process that prioritizes spatially proximate tokens.
We therefore modulate candidate-to-selected similarities by spatial sparsity. For each candidate $i\notin\mathcal{S}$, compute the normalized nearest-selected distance $\bar{\delta}_{i}$, and apply the BSS gain in Eq. (5) with a $\lambda>0$ (we use $\lambda=0.5$). Because $\bar{\delta}_{i}$ grows with spatial distance from $\mathcal{S}$, (5) makes farther candidates appear more similar to some selected token. So they are judged more redundant and thus deferred. This yields a near-to-far centrifugal growth that first fills local details then expands outward.
Given $i\in\mathcal{C}$, the non-duplication score is defined as:
$r_{i}\;=\;1-\max_{j\in\mathcal{S}}\widetilde{M}_{ij},$ (7)
We then sort the candidates by $r_{i}$ in descending order and process them in parallel batches of size $B$, where $\widetilde{M}_{ij}$ is recomputed within each batch. A candidate $i$ will be accepted in loop $t$ if:
$\max_{j\in\mathcal{S}}\widetilde{M}_{ij}\;<\;\tau^{(t)},$ (8)
where, $\tau^{(0)}=0.8,\;\tau^{(t{+}1)}=\tau^{(t)}+\Delta\tau,\;\Delta\tau=0.1.$
Due to the BSS gain, the upper bound of $\widetilde{M}_{ij}$ increases from $1$ to $(1+\lambda)$. The strict-to-loose schedule prevents early admission of isolated, far-away tokens. We iterate loops until $|\mathcal{S}|=R$ or no new tokens are added.
3.3.3 Stage 3: Recovery from the Discarded via SWA
While the centrifugal selection process effectively preserves more complete details, it inevitably discards many outermost tokens that may still contain complementary semantic cues. To mitigate this loss, we introduce a Similarity-Weighted Aggregation (SWA) mechanism that recovers useful information from discarded tokens and reintegrates it into the most relevant retained tokens.
We denote the discarded set as $\mathcal{D}=\{0,\dots,N-1\}\setminus\mathcal{S}$. Each discarded token $u\in\mathcal{D}$ is assigned to its most similar retained token:
$j^{\star}(u)=\arg\max_{j\in\mathcal{S}}M_{uj},$ (9)
which forms clusters $\mathcal{D}_{j}=\{u\in\mathcal{D}\,\And\,j^{\star}(u)=j\}$ for each retained token $j\in\mathcal{S}$. We then compute similarity-normalized weights for each cluster $\mathcal{D}_{j}$ and aggregate the hidden states $\mathbf{H}_{u}$ of the discarded tokens:
$\alpha_{u\to j}$ $=\frac{M_{uj}}{\sum_{u^{\prime}\in\mathcal{D}_{j}}M_{u^{\prime}j}+\varepsilon},$
$\mathbf{E}_{j}$ $=\sum_{u\in\mathcal{D}_{j}}\alpha_{u\to j}\,\mathbf{H}_{u},$ (10)
$\mathbf{H}_{j}$ $=\beta\,\mathbf{H}_{j}+(1-\beta)\,\mathbf{E}_{j},\qquad\beta\in(0,1),$
where $\alpha_{u\to j}$ denotes the normalized aggregation weight between a discarded token $u$ and its assigned retained token $j$, and $\mathbf{E}_{j}$ represents the similarity-weighted hidden state aggregated from all discarded tokens mapped to $j$. The term $\varepsilon=10^{-8}$ ensures numerical stability. We set $\beta=0.3$ in practice, meaning that $70\%$ of the aggregated hidden state $\mathbf{E}_{j}$ is incorporated into the final representation of the retained token.
3.3.4 Complexity
Channel screening and similarity construction cost $O(Nd)$ and $O(N^{2}q)$, respectively. Spatial distances cost $O(N^{2})$ once. Each greedy loop performs batched tensor ops equivalent to $O(N|\mathcal{S}|)$ (indexing and selecting the minimum spatial distance). The number of loops is small in practice due to the annealed threshold. The aggregation stage performs a most similar assignment (using argmax) and a similarity-weighted matrix-vector update for each retained token.
[…]
Upper Bound, 576 Tokens (100%)
LLaVA-1.5-7B 61.95 64.7 1862 85.9 69.46 58.17 297 66.18 57.98 100.0%
[…]
FastV (ECCV24) 57.94 61.2 1829 78.2 68.77 57.37 293 63.60 56.97 96.45%
SparseVLM (ICML25) 59.50 62.5 1782 85.7 68.57 57.70 294 64.23 57.57 97.93%
PDrop (CVPR25) 57.19 63.2 1787 80.3 69.11 56.16 278 59.34 56.01 95.04%
MustDrop (arXiv24) 58.20 62.3 1728 83.6 69.20 55.87 314 62.77 56.57 97.13%
DART (EMNLP25) 60.08 63.6 1855 82.9 69.06 57.52 296 64.62 57.17 98.40%
DivPrune (CVPR25) 60.02 62.0 1771 87.7 68.72 56.64 293 63.89 57.29 97.80%
Ours 61.31 63.5 1788 86.0 69.96 57.97 290 65.34 57.78 98.85%
[…]
FastV (ECCV24) 55.86 56.1 1721 72.2 68.86 56.95 288 61.41 55.87 92.95%
SparseVLM (ICML25) 58.42 60.0 1753 85.5 68.67 56.67 278 63.63 56.78 96.08%
PDrop (CVPR25) 54.43 61.1 1611 73.1 69.05 53.59 268 56.44 53.41 90.34%
MustDrop (arXiv24) 56.19 61.1 1690 79.1 68.50 56.04 306 60.53 55.48 94.80%
DART (EMNLP25) 58.70 63.1 1826 80.7 68.96 56.42 294 63.41 56.58 97.00%
DivPrune (CVPR25) 59.35 61.6 1713 87.5 68.67 55.92 285 62.34 56.56 96.40%
Ours 61.00 62.5 1768 85.4 69.31 57.32 294 64.35 57.18 98.07%
[…]
FastV (ECCV24) 51.66 48.0 1564 59.7 67.91 55.35 249 55.48 52.04 84.60%
SparseVLM (ICML25) 53.75 56.2 1593 78.0 69.71 53.47 203 56.75 53.42 87.61%
PDrop (CVPR25) 41.9 33.3 1092 55.9 68.60 45.90 250 54.73 52.98 75.42%
MustDrop (arXiv24) 52.71 60.0 1533 68.6 68.96 54.49 278 54.68 53.48 89.05%
DART (EMNLP25) 56.38 60.6 1751 74.6 68.67 54.30 271 59.29 55.52 92.71%
DivPrune (CVPR25) 57.74 60.1 1619 86.2 67.97 54.60 271 60.19 55.57 93.68%
Ours 59.15 61.3 1752 82.2 68.37 56.05 279 62.24 56.63 95.61%
Table 1: Comparative experiments on image understanding are performed on LLaVA-1.5-7B.
4 Experiment
[…]
Upper Bound, 576 Tokens (100%)
LLaVA-1.5-13B 63.22 68.6 2826 86.4 72.68 61.19 331 68.23 60.87 100.0%
[…]
FastV (ECCV24) 55.75 55.1 1670 69.5 72.58 56.1 270 59.81 57.19 84.75%
DART (EMNLP25) 57.06 65.7 1736 76.0 73.87 55.84 264 61.76 57.54 88.12%
DivPrune (CVPR25) 57.93 64.3 1754 84.8 71.64 57.38 287 62.31 57.32 90.20%
Ours 61.21 65.7 1779 84.7 72.43 58.74 306 64.91 59.23 92.68%
Dynamic Resolution (Max pixels = 1344 $\times$ 336), Dynamic Height and Width
Upper Bound, Dynamic Tokens, Max Tokens = 576 $\times$ 4 = 2304 (100%)
LLaVA-Next-7B 64.25 67.9 1852 86.5 70.35 64.84 518 70.17 58.58 100.0%
[…]
FastV (ECCV24) 57.49 54.5 1686 76.9 68.22 49.84 264 61.2 55.35 84.02%
DART (EMNLP25) 59.91 63.5 1718 81.7 68.12 55.65 357 63.19 57.16 90.35%
DivPrune (CVPR25) 59.83 63.9 1746 83.6 67.8 49.32 317 65.28 56.46 89.02%
Ours 61.38 64.1 1709 84.3 68.32 56.04 357 66.2 57.37 91.60%
Table 2: Comparative experiments on image understanding are performed on LLaVA-1.5-13B and LLaVA-Next-7B.
[…]
Fixed Resolution (Min pixels = Max pixels = 1280 $\times$ 28 $\times$ 28), Dynamic Height and Width
Upper Bound, About 1350 Tokens (100%)
Qwen2-VL-7B-Instruct 61.64 81.6 2321 88.5 85.42 82.73 796 76.65 50.21 100.0%
[…]
FastV (ECCV24) 60.10 79.6 2319 87.2 82.94 81.28 646 73.52 49.43 96.04%
DART (EMNLP25) 59.55 78.4 2259 86.9 82.25 78.06 651 72.54 49.83 94.95%
DivPrune (CVPR25) 61.11 79.4 2199 88.0 83.39 80.07 704 75.03 49.1 96.57%
Ours 61.54 79.5 2273 87.9 84.28 81.01 750 74.90 49.76 98.02%
[…]
FastV (ECCV24) 58.90 77.6 2290 86.4 82.35 79.43 569 70.95 48.77 93.39%
DART (EMNLP25) 57.64 77.0 2181 86.0 80.47 74.10 586 70.20 48.83 91.69%
DivPrune (CVPR25) 60.59 77.8 2202 87.8 82.10 77.74 607 73.61 48.65 94.11%
Ours 61.49 78.1 2266 88.2 83.49 79.51 702 74.09 49.03 96.57%
[…]
FastV (ECCV24) 55.65 73.5 2174 81.6 78.53 72.86 454 64.79 46.59 86.73%
DART (EMNLP25) 55.86 73.0 2087 81.6 78.14 65.10 481 65.02 46.59 85.60%
DivPrune (CVPR25) 59.17 74.8 2059 86.5 80.22 71.65 472 70.65 47.29 88.93%
Ours 60.49 76.1 2158 87.4 80.81 76.15 581 72.14 48.35 92.58%
Table 3: Comparative experiments on image understanding are performed on Qwen2-VL-7B-Instruct.
4.1 Benchmarks and Base VLMs
To comprehensively evaluate the capability of our proposed VLM-Pruner, which is aimed at multimodal reasoning and understanding, we conduct extensive experiments on a diverse set of benchmark datasets and VLMs with varying architectures. Specifically, we select a broad range of commonly used tasks and benchmarks covering multiple-choice question answering (QA), optical character recognition (OCR), and open-ended QA. The evaluation includes 9 image-language benchmarks, namely GQA [gqa], MMB [mmbench], MME [mme], POPE [pope], SQA [sqa], TextVQA [textvqa], OCRBench [ocrbench], SEEDBench [seed], and OKVQA [okvqa], as well as 4 video-language benchmarks, namely EgoSchema [egoschema], NExTQA [nextqa], VideoMME [videomme], and EgoPlan [egoplan]. For base VLMs, we choose image-based models including LLaVA-1.5-7B [llava15], LLaVA-1.5-13B [llava15], LLaVA-Next-7B [llavanext], and Qwen2-VL-7B [qwen2vl], together with the video-based model LLaVA-Video-7B [llavavideo]. Please refer to the supplementary materials for more details.
4.2 Implementation Details
For LLaVA-1.5 [llava] with a fixed number of visual tokens, we set $H=W=24$ ($N=576$), so that $D_{\max}=24\sqrt{2}$, which exactly matches the normalization in Eq. (5). For LLaVA-Next [llavanext] and Qwen2-VL [qwen2vl], which adopt a dynamic number of tokens depending on the image resolution, $D_{\max}$ is computed according to the actual height and width of the token feature map. For LLaVA-Video [llavavideo], which involves three spatial-temporal dimensions (height, width, and frames), the computation is correspondingly performed in the 3D coordinate space. Meanwhile, we set $\lambda=0.5$, $\tau^{(0)}=0.8$, $\Delta\tau=0.1$, $q=256$, $\beta=0.3$, and $B=16$.
All experiments are performed on a setup consisting of 8 GPUs, each with 32 GB of VRAM. We adhere to the official evaluation protocols established for both LLaVA-1.5 and LLaVA-Next. Additionally, for evaluating Qwen2-VL and LLaVA-Video, we utilize the lmms-evals package [lmmseval]. All results are obtained with a batch size of 1.
4.3 Comparative Experiment
We select methods based on importance-based baselines, such as FastV [fastv], SparseVLM [sparsevlm], PDrop [pyramiddrop], and MustDrop [mustdrop], as well as methods based on redundancy-based baselines, such as DART [dart] and DivPrune [divprune], as the comprehensive baselines for comparison.
4.3.1 Image-Language Understanding
LLaVA-1.5-7B. We evaluate VLM-Pruner under a fixed visual token budget, as shown in Table 1. Across 192, 128, and 64 tokens, our pruning preserves 98.85%, 98.07%, and 95.61% of the upper bound Avg. At 64 tokens with $\downarrow 88.9\%$ pruning, VLM-Pruner ranks first on seven benchmarks and remains competitive under extreme sparsity. The proximity prior in BSS concentrates selections on informative regions and retains fine grained details, which yields a favorable performance and sparsity tradeoff and increasing gains over baselines as the budget tightens.
LLaVA-1.5-13B. We study the larger capacity setting with a fixed visual token budget, as shown in Table 2. At 64 tokens with $\downarrow 88.9\%$ pruning, our approach attains an Avg. of 92.68% and exceeds DivPrune by +2.48%, DART by +4.56%, and FastV by +7.93%. VLM-Pruner also maintains first place on seven benchmarks, which indicates that redundancy and detail balancing transfers to higher capacity VLMs without loss of relative advantage. The relative margins widen on categories that require fine-grained grounding (GQA) and OCR recognition (OCRBench), suggesting that VLM-Pruner better preserves evidence with more comprehensive details than redundancy-based methods.
LLaVA-Next-7B. We examine the dynamic resolution setting where the token count varies with image size, as shown in Table 2. With $\downarrow 88.9\%$ pruning, VLM-Pruner attains an Avg. of 91.60% and ranks first on eight benchmarks. This demonstrates strong adaptability to fluctuations in token count induced by resolution changes, which is essential for maintaining coverage of local evidence as the visual grid varies, thereby reducing performance loss relative to redundancy-based methods.
Qwen2-VL-7B. We evaluate architectural generality across three pruning rates, as reported in Table 3. The Avg. improvement increases with sparsity and reaches +3.65% at $\downarrow 88.9\%$. On OCRBench, which is sensitive to small-size text, VLM-Pruner attains a Final score of 581 versus 481 for DART, an absolute gain of +12.56%. These results indicate that sparsity-aware selection preserves fine-grained visual cues and remains effective across architectures.
Method EgoSch. Acc. NExTQA EM/WUPS VideoMME P-score EgoPlan Acc. Avg.
Max frames = 16
Upper Bound, Max Tokens = 16 $\times$ 182 = 2912 (100%)
LLaVA-Video-7B 55.44 78.75/34.11 59.85 33.80 100%
[…]
FastV (ECCV24) 43.97 67.91/30.66 50.44 28.27 84.96%
DART (EMNLP25) 48.36 72.84/31.66 53.59 28.90 89.84%
DivPrune (CVPR25) 46.27 72.42/31.87 53.96 27.73 88.52%
Ours 47.33 72.93/32.02 54.96 29.58 90.55%
Table 4: Comparative experiments on image understanding are performed on LLaVA-Video-7B-Qwen2.
4.3.2 Video-Language Understanding
LLaVA-Video-7B-Qwen2. We evaluate the video and language capability with $\downarrow 88.9\%$ pruning that retains 20 tokens per frame from the original $13\times 14=182$ tokens, as shown in Table 4. VLM-Pruner surpasses prior work across four benchmarks and attains NExTQA 72.93/32.02, giving absolute improvements of +0.09 EM and +0.15 WUPS over the second best. On VideoMME it reaches a P-score of 54.96, an absolute gain of +1.00. On EgoPlan it achieves an Accuracy of 29.58%, an absolute gain of +0.68%. The overall Avg. is 90.55% versus 89.84% for the second best, an absolute gain of +0.71%, which demonstrates robustness to severe token reduction in spatiotemporal reasoning. BSS computed in 3D coordinates supports stable selection across height, width, and time and avoids dispersed yet incomplete token sets that appear in redundancy-based strategies.
4.4 Efficiency Analysis
We analyze the efficiency of VLM-Pruner on LLaVA-1.5-7B and Qwen2-VL-7B. Specifically, we report FLOPs following [fastv] and total inference time on the POPE and OCRBench benchmarks under a $\downarrow$88.9% pruning rate (i.e., retaining 11.1% of visual tokens), as shown in Table 5. VLM-Pruner outperforms the importance-based baselines SparseVLM and MustDrop in both speedup and FLOPs. In comparisons with redundancy-based methods, DivPrune exhibits slightly fewer FLOPs. This is because it prunes at layer 0, while VLM-Pruner and DART start at layer 2. However, this DivPrune setting comes at the cost of significantly poorer performance compared to our method (Sec. 4.3). VLM-Pruner’s inference speed matches DART. On Qwen2-VL-7B, which has more raw visual tokens, VLM-Pruner excels in detail-intensive tasks like OCRBench. It achieves the fastest speed (1.60×) and the best performance (Table 3). These results demonstrate VLM-Pruner’s superior capability for handling high-resolution inputs.
[…]
POPE OCRBench
LLaVA-1.5-7B 576 33:21 7:14 100%
[…]
Qwen2-VL-7B 100% 11:33:16 1:28:16 100%
[…]
Table 5: Inference costs of the number of tokens, Total-Time, Speedup, and FLOPs.
4.5 Ablation Study
Structural decomposition analysis (Table 6).
(i) Stage 1. Replacing the max-min diverse pivot initialization with a direct Top-4 L1-distance keys seeding the Avg. by -1.27%, indicating that semantically diverse pivots cover the target region better than simply selecting large-magnitude key vectors. (ii) Stage 2. Substituting the selector of stage 2 with DART or DivPrune reduces the Avg. to 90.46% and 46.47%, respectively, showing clear degradation because pure redundancy removal disperses coverage and loses detail. Removing the normalized nearest-distance term in BSS (i.e., omitting $\bar{\delta}_{i}(S)$) yields a drop of -1.11%, which underscoring the benefit of centrifugal expansion for preserving details. (iii) Stage 3. Dropping similarity-weighted aggregation lowers the Avg. from 95.30% to 95.07%. Although the gain is modest, it is consistent and improves performance without increasing the token count, so it is complementary to the first two stages.
Ablation studies on hyperparameters (Fig. 4).
(a) Four pivots are most robust. Fewer fail to cover multiple targets. More tend to over-cover edges. (b) Variance-based channel selection is best around $q=256$. Larger $q$ adds redundancy and complexity. (c) The initial threshold $\tau^{(0)}=0.8$ performs best. A smaller threshold admits far candidates too early and causes dispersion. A larger threshold increases iterations and hinders effective expansion. (d) Increasing token batch $B$ reduces latency while performance remains stable. $B=16$ offers a good speed-performance trade-off.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
Ours 1752 82.2 56.05 279 62.24 56.63 95.30%
[…]
Top-4 keys 1754 80.8 55.61 266 61.90 56.37 94.03%
[…]
+ DART 1727 75.8 53.15 248 60.41 55.42 90.46%
+ DivPrune 1000 28.8 42.5 24 37.50 31.17 46.47%
w/o $\bar{\delta}_{i}(S)$ 1726 81.9 55.54 270 61.82 56.41 94.19%
[…]
w/o stage 3 1795 81.2 55.72 275 61.91 56.56 95.07%
Table 6: Structural decomposition analysis in three stages.
Figure 4: Ablation studies on hyperparameters on LLaVA-1.5-7B. (a) Number of pivots $\kappa$, (b) Top-$q$ highest variance channels, (c) $\tau^{(0)}$, (d) Total inference time (affected by token batch size $B$).
5 Conclusion
We present VLM-Pruner, a training-free Centrifugal token pruning paradigm that explicitly balances redundancy and spatial sparsity to preserve fine-grained visual details while aggressively reducing compute. By combining a BSS criterion with parallel greedy selection and similarity-weighted aggregation, our approach attains strong end-to-end speedups without sacrificing understanding quality. Extensive experiments across 5 VLMs and 13 image/video benchmarks show consistent state-of-the-art performance under severe pruning, with particular robustness on detail-sensitive tasks such as grounding and OCR. We believe these establish an effective foundation for the practical and high-efficiency multimodal reasoning.
[…]
Supplementary Material
6 Additional Experiments
To further validate the effectiveness of token pruning by VLM-Pruner and the reasonableness of the hyperparameter settings, we present several additional experimental results on LLaVA-1.5-7B with the pruning rate of 88.9%, focusing on (a) number of pivots $\kappa$, (b) top-$q$ highest variance channels, (c) threshold $\tau^{(0)}$ of token selection, (d) token batch size $B$, (e) aggregation weight $\beta$, and (f) pruning layer $i$.
6.1 Ablation Study on the Number of Pivots $\kappa$
In the experiment on the number of pivots as shown in Table 7, we observe that the performance varied with different values of $\kappa$. The optimal performance was achieved with $\kappa=4$, resulting in an average performance of 95.30%. This suggests that four pivots strike the best balance, providing sufficient coverage of the target regions without overfitting to a specific area. Fewer pivots (e.g., $\kappa=1$ or $\kappa=2$) fail to capture enough semantic diversity, while too many pivots (e.g., $\kappa=7$ or $\kappa=8$) tend to overfit and capture redundant tokens from less important regions.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
1 1737 81.6 55.93 271 61.83 56.29 94.36%
2 1729 81.9 55.66 271 62.33 56.52 94.46%
3 1725 82.4 55.80 269 62.36 56.33 94.41%
4 1752 82.2 56.05 279 62.24 56.63 95.30%
5 1734 81.6 55.97 273 62.21 56.72 94.68%
6 1732 82.0 56.11 281 62.17 56.81 95.25%
7 1722 81.5 55.88 274 62.11 56.67 94.55%
8 1707 82.0 56.03 273 62.13 56.33 94.40%
Table 7: Ablation study on the number of pivots $\kappa$.
6.2 Ablation Study on the Top-$q$ Highest Variance Channels
For the experiment on selecting the top-$q$ highest variance channels as shown in Table 8, we find that setting $q=256$ achieved the best performance, with an average score of 95.30%. This configuration retains enough channels to preserve essential visual information while minimizing redundancy. Larger values of $q$ result in diminishing returns, as they introduce more redundant information and computational complexity, which doesn’t contribute significantly to performance improvement. On the other hand, smaller values of $q$ lead to the loss of critical visual features, reducing the overall model performance. Therefore, $q=256$ is the optimal setting, offering a balanced trade-off between computational efficiency and model performance.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
128 1765 81.8 56.25 277 62.02 56.70 95.25%
256 1752 82.2 56.05 279 62.24 56.63 95.30%
512 1738 81.3 55.87 270 62.21 56.59 94.42%
1024 1743 81.1 55.51 268 62.26 56.71 94.26%
4096 (raw) 1752 81.9 55.97 279 62.18 56.41 95.14%
Table 8: Ablation study on the Top-$q$ highest variance channels.
6.3 Ablation Study on the Threshold $\tau^{(0)}$ of Token Selection
In the experiment on the token selection threshold $\tau^{(0)}$ as shown in Table 9, we observe that setting $\tau^{(0)}=0.8$ resulted in the best overall performance of 95.30%. This threshold ensures that the model prioritizes selecting tokens that are spatially close to previously selected ones, effectively avoiding the early inclusion of distant or irrelevant tokens. If the threshold is set too low, distant tokens are selected too early, causing the token set to become scattered and less effective in representing fine-grained details. Conversely, setting the threshold too high causes an overemphasis on spatial distance, resulting in a more compact token selection that introduces increased redundancy. Therefore, a threshold of 0.8 strikes a balance between selecting non-redundant tokens and preserving finer-grained details.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
0.6 1753 81.9 55.00 276 61.82 56.48 94.64%
0.7 1704 82.5 55.68 276 61.93 56.71 94.60%
0.8 1752 82.2 56.05 279 62.24 56.63 95.30%
0.9 1719 78.3 54.71 269 61.29 56.39 92.99%
1.0 1725 73.9 53.66 265 59.87 55.31 91.00%
Table 9: Ablation study on the threshold $\tau^{(0)}$ of token selection.
6.4 Ablation Study on the Token Batch Size $B$
In the experiment on token batch size $B$ as shown in Table 10, the results show that a batch size of $B=16$ provides the best performance and fastest inference speed, with an average score of 95.30% and a total time of 91 minutes. Larger batch sizes, such as $B=32$, may reduce inference time, but they may lead to the failure of the BSS criterion due to fewer iterations. Smaller batch sizes result in longer inference times, even slower than the initial LLaVA-1.5-7B at 131 minutes. The token batch size of $B=16$ offers an optimal trade-off, minimizing latency while maintaining the quality of the selected tokens.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 131 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
1 160 1755 83.6 56.06 275 62.95 56.88 95.62%
2 124 1742 83.5 55.96 279 62.73 56.93 95.64%
4 106 1730 83.3 56.11 278 62.57 56.70 95.38%
8 96 1726 83.5 56.26 274 62.42 56.82 95.20%
16 91 1752 82.2 56.05 279 62.24 56.63 95.30%
Table 10: Ablation study on the token batch size $B$.
6.5 Ablation Study on the Aggregation Weight $\beta$
In the aggregation weight $\beta$ experiment, as shown in Table 11, we focus on minimizing the influence of Stage 3 to emphasize the importance of Stage 1 and 2, which manage buffering for spatial sparsity and token selection, while ensuring the preservation of fine-grained details. The results show that $\beta=0.3$ yields an average score of 95.30%, providing a moderate improvement over the 95.07% achieved without Stage 3. It also enables effective aggregation of discarded token information without overemphasizing it. This setting further preserves fine-grained details, as demonstrated by the best score of 279 achieved on OCRBench.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
0.2 1737 82.1 55.86 278 62.26 56.41 94.98%
0.3 1752 82.2 56.05 279 62.24 56.63 95.30%
0.4 1754 82.4 56.26 277 62.28 56.72 95.34%
0.5 1773 82.4 56.44 278 62.36 56.80 95.66%
0.6 1790 82.0 56.16 278 62.36 56.87 95.68%
0.7 1799 81.5 56.23 277 62.33 56.80 95.60%
0.8 1808 81.5 56.07 275 62.14 56.81 95.47%
w/o stage 3 1795 81.2 55.72 275 61.91 56.56 95.07%
Table 11: Ablation study on the aggregation weight $\beta$.
To further emphasize the core innovation of VLM-Pruner, the BSS criterion, we validate the model performance on Qwen-VL-7B after removing Stage 3, as shown in Table 12. The results show that even without Stage 3, our method still outperforms existing baselines by an absolute margin of approximately 5%. This highlights the effectiveness of the proposed centrifugal token pruning paradigm.
[…]
Qwen2-VL-7B Upper Bound (100%)
Vanilla 2321 88.55 82.73 796 76.65 50.21 100.0%
[…]
w/o Stage 3 2197 86.4 74.89 561 70.69 48.20 90.25%
FastV 2174 81.6 72.86 454 64.79 46.59 84.72%
DART 2087 81.6 65.10 481 65.02 46.59 83.14%
DivPrune 2059 86.5 71.65 472 70.75 47.29 86.47%
Ours 2158 87.4 76.15 581 72.14 48.35 91.20%
Table 12: Additional structural decomposition analysis.
6.6 Ablation Study on the Pruning Layer $i$
In the layer-wise pruning experiment summarized in Table 13, pruning at layer 2 achieves the best balance and an average performance of 95.30% by successfully preserving fine-grained visual details while reducing task-irrelevant redundancy. Unlike DivPrune, VLM-Pruner cannot prune layer 0 because it relies on token keyskeys from the previous layer to perform coarse semantic abstraction. Pruning at layer 1 leads to notable performance degradation, likely because the text and visual tokens have not yet fully interacted, which hampers the model’s ability to focus on prompt-relevant regions and results in suboptimal token selection. Conversely, pruning at later layers such as layer 3 or 4 may remove more less important details, and this also comes at the cost of lower model performance and increased computational overhead in FLOPs.
[…]
LLaVA-1.5-7B Upper Bound, 576 Tokens (100%)
Vanilla 1862 85.9 58.17 297 66.18 57.98 100.0%
[…]
1 1689 82.0 55.68 267 61.54 56.30 93.65%
2 1752 82.2 56.05 279 62.24 56.63 95.30%
3 1725 82.9 55.85 274 62.39 56.53 94.86%
4 1759 83.3 55.70 270 62.18 56.48 94.91%
Table 13: Ablation study on the pruning layer $i$.
7 Details for Benchmarks
We select a broad range of commonly used tasks and benchmarks. Specifically, the evaluation includes 9 image-language benchmarks, namely GQA [gqa], MMB [mmbench], MME [mme], POPE [pope], SQA [sqa], TextVQA [textvqa], OCRBench [ocrbench], SEEDBench [seed], and OKVQA [okvqa], as well as 4 video-language benchmarks, namely EgoSchema [egoschema], NExTQA [nextqa], VideoMME [videomme], and EgoPlan [egoplan].
7.1 Image-Language Benchmarks
GQA
GQA is designed to evaluate visual scene understanding and compositional reasoning capabilities of VLMs on real-world imagery.
MMB
MMBench provides a hierarchical evaluation framework for VLMs, organizing 20 distinct capability dimensions into three progressive levels (L1 to L3).
MME
MME assesses both perceptual and cognitive abilities of VLMs, where perception includes optical character recognition (OCR), coarse- and fine-grained object recognition, and cognition encompasses commonsense reasoning, numerical calculation, text translation, and code reasoning.
POPE
POPE quantifies object hallucination in VLMs by reformulating the problem as a binary judgment task on the presence of objects in an image.
SQA
ScienceQA measures scientific reasoning proficiency across natural science, social science, and language science domains, supplemented with detailed explanations and lectures.
TextVQA
TextVQA evaluates the integration of visual and textual information, requiring models to read and reason about text present in images.
OCRBench
OCRBench provides a comprehensive evaluation of text-related capabilities, covering representative tasks such as scene text recognition, document-oriented understanding, key information extraction, and handwritten mathematical expression recognition.
SEEDBench
SEEDBench assesses multimodal understanding across text and visual modalities, including scene classification, object detection, and attribute recognition.
OKVQA
OKVQA examines the ability of VLMs to leverage external knowledge, as the image content alone is insufficient to answer the questions.
7.2 Video-Language Benchmarks
EgoSchema
EgoSchema focuses on long-term video-language understanding, requiring capabilities such as scene comprehension, object state tracking, and extended visual memory.
NExTQA
NExTQA challenges VLMs to interpret temporal and causal structures in videos, emphasizing reasoning about action causality, temporal progression, and object interactions.
VideoMME
VideoMME is the first benchmark dedicated to holistic video analysis, systematically evaluating model performance across varying video durations (short, medium, long) and input modalities (video frames, subtitles, audio).
EgoPlan
EgoPlan assesses VLMs as embodied task planners, involving interactions with hundreds of objects under complex visual settings to evaluate the prediction of feasible actions.